OpenStudy (anonymous):

A student traveled 239km in 12 days. Qt the same ratio, how far would the student travel in 48days? ____km

OpenStudy (anonymous):

This is an ratio and proportion problem: 239km/12day =xkm/48day. You can cross multiply and get the answer for x. You will get 956km. Or you can think 48 days is 4 times 12 days. if a student traveled 239 km in 12days, s/he must travel 4-239km in 48 days

Directrix (directrix):

239 / 12 = x / 48 where x is the number of km for 48 days 12 x = (239)*(48) / 12 x = 956 km Check my work, please.

OpenStudy (anonymous):

correct!

OpenStudy (anonymous):

so 956km